OpenStudy (anonymous):

A)The cube roots of 1 + i B)Give the exact roots. (Express θ in radians.)

OpenStudy (tkhunny):

Really, @cebroski ? You will need DeMoivre, right? Show it and demonstrate it.

OpenStudy (anonymous):

@tkhunny By DeMoivre, \[Z ^{N} = R ^{N}( \cos(N \theta) + i \sin(N \theta))\]

OpenStudy (anonymous):

where Z is a complex number, R is the modulus, N is a rational number, sin and cos are trigonometric functions and i = sqrt(-1)

OpenStudy (anonymous):

Let Z = 1 + i, N = 1, so that we may solve for R and theta. \[Z^1 = 1 + i = R^1(\cos \theta + i \sin \theta)\] For what values of theta and R make the above equation true? \[\theta = \frac{ -7\pi }{ 4 }, \frac{ \pi }{ 4 }, \frac{ 9\pi }{ 4 }\] \[R = \sqrt {2}\]

OpenStudy (anonymous):

Now, let us find the cube root of 1 + i. Let Z = 1 + i, N = 1/3, \[R = \sqrt {2}\] \[\theta = \frac{ -7\pi }{ 4 }, \frac{ \pi }{ 4 }, \frac{ 9\pi }{ 4 }\]

OpenStudy (anonymous):